OpenStudy (anonymous):

Write an expression in simplest form for the area of a square with a side of 8x3 units

OpenStudy (anonymous):

8x^ 3 4 units is the side of the square Area of square = side x side Area = 8x^3 * 8 x^3 = 8 *8 x^3 * x^3 = 64 * x^(3+3) = 64x^6